OpenStudy (anonymous):

solve 3x^2+4x+5=0 a.-10+-i√22/3 b.-2+-i√11/3 c.-2+-i√22/12 d.-2+-5i√11/3

OpenStudy (mathmate):

b.(-2+-i√11)/3

OpenStudy (anonymous):

can you show me the steps please?

OpenStudy (mathmate):

3x^2+4x+5=0 Use the quadratic equation: a=3, b=4, c=5 x=(-b+-sqrt(b^2-4ac))/2a =(-4+-sqrt(4^2-4*3*5))/(2*3) =(-4+-sqrt(44))/(6) =(-2+-sqrt(11))/3